OpenStudy (anonymous):

A music store sold 1,112 music CDs one month. This was 20% less than the number of music CDs the store sold the previous month. The following equation represents this situation, where x represents the number of music CDs sold the previous month. x − 0.2x = 1,112 How many music CDs did the store sell in both months combined?

OpenStudy (anonymous):

Numbers of CDs sold previous month = x 20% less than x means 80% of x = 1112 0.8 x = 1112 x = 1112/0.8 = 1390 CD's sold previous month is 1390

OpenStudy (anonymous):

Yes x -0.2x = 1112 represents the number of music CDs sold previous month. Total CD's sold is 1112 + 1390 =2502

OpenStudy (anonymous):

20% lessthan CDs sold in previous month therefore consider tht 1112 is the 80 percent of no. of cds sold therefore 80%of x=1112=1390Cds sold in perv month x − 0.2x = 1,112 1390-278=1112 1112+1390=2052